OpenStudy (anonymous):

what is the partial derivative p=4x^3y^3

OpenStudy (anonymous):

\[p=4x ^{3}y^{3}\]

OpenStudy (anonymous):

partial derivative of p /partial derivative of y this is written in my book i have to take out this...

OpenStudy (lgbasallote):

you mean dp/dy?

OpenStudy (lgbasallote):

i mean that curved d thingy...you know what i mean :))

OpenStudy (lalaly):

\[\large{\frac{\partial p}{\partial y}}=3 \times 4x^3y^2=12x^3y^2\]